2.80 moles of an ideal gas with CV,m=3R/2 undergoes the transformations described in the following list from an initial state described by T = 310. K and P = 1.00 bar.

The gas is heated to 615 K at a constant volume corresponding to the initial volume. Calculate q for this process
Answer #1

Aus) given = 2.8 mal Cu = BR Ti = 310k T2 = 615k AT = T2-T, = 615 - 310 = 305k Solution q - DE + w. -090 Where DE= nCVOT = mx